Help, I just can't do it!
x is 3 years older than y
z is twice as old as X
4 years ago, x was twice as old as y
How old is x ?

Answer by stanbon(75887)

x is 3 years older than y
z is twice as old as X
4 years ago, x was twice as old as y
How old is x ?
---------------------
When you have three variables, you often need three equations:
x = y+3
z = 2x
(x-4) = 2(y-4)
----------------
Rearrange:
x - y + 0 = 3
2x + 0 - z = 0
x -2y + 0 = -4
----------------------
Solve by any method to get:
x = 10
y = 7
z = 20
==================
Cheers,
Stan H.